PLEASE PROVIDE STEP BY STEP AS WELL AS FINAL ANSWERQuestion 7 (Mandatory) (2 points) 7) Use the given data to find the equation of the regression line. Round the final values t
Answer #1

Solution:

From given data , we prepare a table.

X Y XY X^2 Y^2
0 8 0 0 64
3 2 6 9 4
4 6 24 16 36
5 9 45 25 81
12 12 144 144 144
n 5
sum(XY) 219.00
sum(X) 24.00
sum(Y) 37.00
sum(X^2) 194.00
sum(Y^2) 329.00
Numerator 207.00
Denominator 329.76
r 0.6277
r square 0.3940
Xbar(mean) 4.8000
Ybar(mean) 7.4000
SD(X) 3.9699
SD(Y) 3.3226
b 0.5254
a 4.8782

Now ,

Slope of the regression line is

b = ΠΣxy - (Σκ) Σy) ΠΣχ2 - (Σx)2

  \therefore b = 0.525

Now , y intercept of the line is

a = Σy - bΣΧ n

  \therefore a = 4.878

The equation of the regression line is

\hat y = a + bx

Answer:  \hat y = 4.878 + 0.525x